What is f(g(2))
if f(x)=3x+4
g(x)=(x+1)^(2)
And can you show the steps as well please

Answer by lwsshak3(11628)

What is f(g(2))
if f(x)=3x+4
g(x)=(x+1)^(2)
g(2)=(2+1)^2=3^2=9
f(x)=3x+4
f[g(2)]=3*9+4=27+4=31